THE REVENUE FOR A SANDWICH SHOP IS DIRECTLY PROPORTIONAL TO ITS ADVERTISING BUDGET. WHEN THE OWNER SPENT $2,000 A MONTH ON ADVERTISING, THE REVENUE WAS $120,000. IF THE REVENUE IS NOW $180,000, HOW MUCH IS THE OWNER SPENDING ON ADVERTISING.
THANKS
JMS

Answer by checkley71(8403)

2000/120000=x/180000 now cross multiply
120000x=2000*180000
120000x=360,000,000
x=360,000,000/120,000
x=$3,000 amount of the advertisement to raise the revenue to $180,000